OpenStudy (anonymous):

Please help me graph the relation and find the domain and the range. [-1, 1/2] [-1/2, -1] [3/2, 0] [2, 3/2]

OpenStudy (anonymous):

\[(-1 , \frac{ 1 }{ 2 } ) , (- \frac{ 1 }{ 2 } , -1) , ( \frac{ 3 }{ 2 } , -1) , (2 , \frac{ 3 }{ 2 })\]

OpenStudy (anonymous):

A.

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

I think you might have the wrong pictures attached because 0 is not in the domain or range

OpenStudy (anonymous):

Those were the ones I'm supposed to graph.

OpenStudy (anonymous):

but at the top of this page you have to sets of numbers which ones go with the pictures

OpenStudy (anonymous):

Is it really that simple and I missed it? TT,TT

OpenStudy (anonymous):

Thank you!